# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v2

@@ Abstract.
Conventional wisdom holds that an efficient interface between an OS running on a CPU and a high-bandwidth I/O device should be based on Direct Memory Access (DMA), descriptor rings, and interrupts: DMA offloads transfers from the CPU, descriptor rings provide buffering and queuing, and interrupts facilitate asynchronous interaction between cores and device with a lightweight notification mechanism.
In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers. Like others before us, we argue that the assumptions that led to this model are obsolete, and in many use-cases use of Programmed I/O (PIO), where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
However, unlike others to date, we push this idea further and show, in a real implementation, the gains in average and tail latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device. We show this using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using PIO over PCI Express (PCIe).
@@ 1. Introduction
Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators are designed to optimize throughput. Built over a PCIe interconnect, descriptor rings in main memory hold queues of operations and completions, and the device performs DMA transactions to read and write both descriptors and payload data. This design, along with the underlying PCIe interconnect, trades off latency for small transactions in favor of high throughput for large ones.
@@
In this paper we question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled hardware accelerators for irregular workloads, or data center Remote Procedure Call (RPC) using small messages. We do this in the context of emerging cache-coherent peripheral interconnects like CXL.cache 3.0 (CXL Consortium, 2024).
@@
The continuing increase in PCIe bandwidth means that, for a given transfer size, the significance of PCIe latency (time to first byte) has become more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 7.
@@
However, we propose a more radical approach, aimed at use-cases where reduced latency is more important than higher maximum throughput: using PIO to Memory-Mapped I/O (MMIO) devices directly, and avoiding DMA, descriptor queues, operation buffering, and interrupts entirely.
@@
We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
@@
We show that, for small (around 1KiB or less) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency penalty. However, a full cache coherence protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe.
Crucially, not only is median latency considerably lower, but the tail of the latency curve is reduced significantly when using PIO over PCIe, and completely eliminated using the cache coherence protocol.
@@
We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices which can interact with an unmodified CPU at the coherence message level. By relaxing traditional coherence protocol assumptions (for example, the independence of cache lines), we can achieve dramatically more efficient communication between a CPU and a device.
@@
In the following section, we provide background and motivation for this work, including the factors that led to the descriptor-based DMA model of I/O, and why it is time to question the underlying assumptions in that model. In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers. We show how to use PIO with PCIe in Section 4.
@@
In Section 5, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol. In Section 5, we go on to present the implementation of some of these protocols on our experimental platform.
@@
Section 6 then compares a traditional DMA-based approach with PIO over PCIe and our new coherence-based protocols using three different use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of a distributed dataflow stream processor to an FPGA. We survey related work in Section 7 and conclude with Section 8.
@@ 2. Background and Motivation
In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space. Both data transfer and control of the device involved reads and writes by the CPU to these registers. This model is also known as memory-mapped I/O (MMIO).
This model made sense when devices were extremely simple, particularly when augmented by the use of device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete. However, the model still requires a synchronous rendezvous between the CPU and device, and moreover requires all data to be transferred via CPU registers (since the CPU is copying data between device registers and main memory).
DMA removes this limitation by allowing the device itself or a dedicated DMA controller to copy the data while the CPU continues to execute other code, providing parallelism and partly decoupling the CPU and device. Further decoupling is provided by descriptor rings: producer/consumer buffers of I/O requests in memory represented as a queue of operations, and read and written by both the CPU and the DMA-capable device. In this case, interrupts need only be used when the queue of descriptors becomes full or empty.
Almost every high-speed device today uses essentially this technique, with minor variations such as how the descriptor queues are formatted, where they are stored, and whether the head and tail of the queue is identified by registers or flags in the descriptors themselves. This model is optimized for throughput – it is most efficient when a large volume of data must be transferred. On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks, up to 64GiB/s for PCIe 5.0 x16.
@@
For the large batch transfers used in GPU-derived programming models (e.g., machine learning workloads), this latency penalty is insignificant when amortized over a large transfer. Indeed, such software has arguably adapted to this model imposed by the hardware (Hooker, 2021).
Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024). The more far-reaching of these proposals are fully symmetric, i.e. devices are first-class participants in a distributed directory-based cache coherence protocol.
@@
All these workloads incur a significant latency penalty when using DMA for data transfer. Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads.
For example, CPU execution and I/O were decoupled using descriptor rings because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
This is no longer true. Server CPUs now have 10s or 100s of cores, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019). Most of the CPU cores have nothing to do except handle small messages coming from the network.
@@
These reasons led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware. We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications?
Note that we are not proposing abandoning the traditional model, or some of its enhancements we survey in Section 7, for throughput-oriented applications, which can continue to ride the current hardware curve for bandwidth. Instead, we argue it is time to examine complementary alternatives.
@@ 3. Experimental platform
Figure 1. XDMA invocation latency comparison.
Figure 2. PIO invocation latency comparison.
@@
Both nodes have PCIe interfaces: the ThunderX-1 has a PCIe Gen.3 x8 interface, while the FPGA has a PCIe Gen.3 x16 interface. In this work we connect these two interfaces with a loopback cable; this essentially replicates the case of an PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
@@
We therefore calibrate against a modern PC for both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 6.1. The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
@@
We see that XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much insignificant. Single transaction latency is almost constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
The performance difference between the machines is due to several factors. For small transfers, the CPU overhead of descriptor setup dominates, and an x86 core is simply faster than a ThunderX-1 core. In addition, the memory system is somewhat faster on the PC. The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
@@
In the second calibration experiment the CPU performs PIO reads and writes over PCIe to the FPGA. We experimented with several different optimizations in this experiment (see Section 4), in particular the use of write combining and vector instructions. We present the best figures obtainable on each platform.
Figure 2 shows that for transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA. Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms). Writes, in contrast, are combined and pipelined at the PCIe interface.
4. Programmed I/O over PCIe
PCIe is complex, and we spent considerable time optimizing CPU-initiated reads and writes for different platforms. We describe here how we arrived at optimally-efficient PIO latency to the FPGA over PCIe.
We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3).
@@ 4. Programmed I/O over PCIe
Figure 3. An invocation using PIO & PCIe. Error handling omitted.
@@
x86 has more optimization opportunities for PIO reads and writes. The Intel Data Streaming Accelerator (Intel, 2024) offers MOVDIR64B and ENQCMD instructions, allowing 64-byte PIO writes with write-combining mappings and user-space descriptor submission without driver intervention. Mapping memory with write-through attributes, as suggested by Tide (Humphries et al., 2024), allows cache line-sized PIO reads. These architecture-specific extensions are not available on ThunderX-1.
@@ 5. PIO over a coherent interconnect
We now turn to programmed I/O over a cache coherent interconnect. A naive approach would have the CPU simply copy data over a coherent interconnect as in conventional PIO. This requires careful cache management on the CPU using explicit cache flush and invalidate instructions, but is appropriate when the hardware standard mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
However, we are more interested in how a coherent interconnect can best perform low-latency, efficient data movement between CPU registers and device hardware, including cases where the conventional notion of the cache as a view over memory contents is not appropriate, such as data arriving over a network.
Scoping:
We scope the scenario further to the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory. Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
We also assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines (although it might not actually cache data itself). Moreover, we take the perspective of the device designer: the device sees individual cache messages, rather than transparently coherent memory.
These assumptions are rare in hardware today but they do hold for proposed standards like CXL 3.0 (CXL Consortium, 2024), and also for the Enzian platform we use in experiments.
Baseline: simple coherence
We first discuss using simple cache coherence to re-apply existing techniques inter-core communication in software, such as FastForward (Giacomoni et al., 2008). Here (Figure 4), the sender (CPU) writes data to a cache line on which the receiver is spinning reading. Assume the line starts in Shared state in both caches. When the sender writes the line (typically atomically via the core’s write buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
Figure 4. Simple inter-core communication over coherence
This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in minimal interconnect traffic (since the receiver spins on a local copy of the line), although the CPU itself must spin. Indeed, this is the technique recently used in CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket.
However, this design assumes that each endpoint can only manipulate its local cache using software instructions (loads, stores, etc.) and also that flush or invalidate instructions should be avoided for performance reasons. This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
@@
A device on a symmetric coherent interconnect can see more cache state and events, and also initiate more operations, than a CPU core can. Note that this does not entail a cache per se on the device at all, indeed this is rarely appropriate.
First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid). It is free to react internally to these events as it chooses, as long as it does not violate the protocol.
Second, the device can issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
@@
Put simply, requests to specific cache lines can be used to signal particular higher-level operations to the device, much as register reads and writes to traditional devices frequently have semantics very different from memory access.
This is an oversimplification, and we discuss the subtleties that arise below. First, however, we present four related communication protocols, layered above MESI-like coherence, which provide for efficient, blocking communication between a conventional CPU and a coherence-aware device.
(a) CPU-initiated, device-homed
(b) Device-initiated and homed
(c) CPU-initiated, CPU-homed
(d) Dev.-initiated, CPU-homed
@@
Communication protocol design variants
We present 4 different but similar communication protocols (shown in Figure 5) which vary based on whether (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device. All provide a synchronous invocation: a transfer of a single cache-line-sized message, followed by a single cache-line response. All use 2 cache lines, which we designate A and B. These protocols can be generalized to larger payloads, and latency will grow linearly with the payload size. Interestingly, they can be made slightly more efficient by using more cache lines, but quickly become complex. We focus on the simple 2-line case in this paper.
The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated invocation on the accelerator using addresses homed on the device (5(a)). An implementation of this protocol is used for the evaluation in Section 6.1. Note that, unlike RPC, invocations are local to the machine and avoid marshaling or unmarshaling of messages.
@@
    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid (we show how this can be conveniently achieved below). To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified. This happens without communication with the device, since B was previously in Exclusive.
@@
    To signal to the device that line B holds a new request, the CPU then executes a load to line A. Since A is Invalid, the cache requests A in Shared from the device. This violates the typical convention that cache line operations are independent, but the device is free to interpret this request for A as a signal from the CPU that B holds a request.
@@
    The device does not respond immediately to the request for A, thereby stalling the load and blocking the requesting core. Instead, it requests B from the CPU’s cache, in Exclusive. The interconnect therefore transfers the request data to the device, and line B duly transitions to Invalid in the CPU’s cache.
@@
    The device reads the data it has received in B, and computes a result on it, which is to be returned to the CPU. It returns this using the response to the CPU’s request for cache line A.

### v3

@@ Abstract.
Conventional wisdom holds that an efficient interface between an OS running on a CPU and a high-bandwidth I/O device should use Direct Memory Access (DMA) to offload data transfer, descriptor rings for buffering and queuing, and interrupts for asynchrony between cores and device.
In this paper we question this wisdom in the light of two trends: modern and emerging cache-coherent interconnects like CXL3.0, and workloads, particularly microservices and serverless computing. Like some others before us, we argue that the assumptions of the DMA-based model are obsolete, and in many use-cases programmed I/O, where the CPU explicitly transfers data and control information to and from a device via loads and stores, delivers a more efficient system.
However, we push this idea much further. We show, in a real hardware implementation, the gains in latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device, and that throughput is competitive with DMA over modern interconnects. We also demonstrate three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using memory-mapped programmed I/O over PCIe.
@@ 1. Introduction
Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), Graphics Processing Units (GPUs), etc. are designed to optimize throughput for large transfers. Built over a PCI Express (PCIe) interconnect, descriptor rings in memory hold queues of requests and completions, and the device uses DMA to read and write both descriptors and payload data. This design, along with the underlying PCIe interconnect, trades off latency for small transactions in favor of high throughput for large ones.
@@
We question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled accelerators for irregular workloads, or Remote Procedure Call (RPC) using small messages. We do this in the context of emerging cache-coherent peripheral interconnects like CXL.mem 3.0 (CXL Consortium, 2024).
@@
The continuing rise in PCIe bandwidth means, for a given transfer size, the significance of PCIe latency (time to first byte) is now more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 6).
@@
We propose a more radical approach, aimed at use-cases where low latency is more important than high maximum throughput: using Programmed I/O (PIO) to issue loads and stores to Memory-Mapped I/O (MMIO) registers directly, avoiding DMA, descriptors, queues, and interrupts entirely.
@@
Using a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA, we investigate the trade-offs between DMA with descriptors, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
@@
We show that, for small ($\leq$ 1KiB) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are much less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency, in line with other recent critiques of DMA (Liew et al., 2025; Sadok et al., 2023; Schuh et al., 2024).
@@
We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices interacting with an unmodified CPU using coherence messages. By relaxing traditional coherence protocol assumptions (for example, the independence of cache lines), we can achieve dramatically more efficient communication between a CPU and a device. A coherence-based message protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe and completely eliminates tail latency, while delivering the same or better throughput.
@@
We first provide background and motivation, including the origins of descriptor-based DMA, and why it is time to question these underlying assumptions. In Section 3 we describe and motivate the experimental hardware platform we use for evaluation, calibrating performance against conventional PC servers. We also decribe PIO with PCIe here.
@@
In Section 4, we discuss why true cache-coherent device interconnects like CXL3.0 are fundamentally different from today’s PCIe, and present a family of protocols that pass messages with low latency between software on an unmodified CPU, and a smart device with message-level access to the coherence protocol. We then present the implementation of these protocols on the hardware platform.
@@
Section 5 compares traditional DMA-based I/O with PIO over PCIe and our new coherence-based protocols using three use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of stream processor operators. We survey related work in Section 6 and conclude with Section 7.
@@ 2. Background and Motivation
In early machines, the CPU interacted with peripherals exclusively using Programmed I/O (PIO): the device exposed MMIO registers, and the processor issued loads and stores to directly to these registers to perform both data transfer and control of the device.
This model made sense when devices were extremely simple, particularly combined with device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete. However, it still requires synchronous rendezvous between CPU and device, and moreover requires all data to be transferred via CPU registers (since the CPU is copying data between device registers and main memory).
DMA removed this limitation by allowing the device itself or a dedicated DMA controller to copy data while the CPU continues to execute other code, providing parallelism and partly decoupling CPU and device. Further decoupling is provided by descriptor rings: producer/consumer buffers of I/O requests in memory, read and written by both CPU and DMA-capable device. In this case, interrupts need only be used when the queue of descriptors becomes full or empty.
Almost every high-speed device today uses this technique, with minor variations such as the format of descriptor queues, where they are stored, and whether the queue head and tail are identified by registers or flags in the descriptors. This model is optimized for throughput – it is most efficient when a large volume of data must be transferred. On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks, up to 64GiB/s for PCIe 5.0 x16.
@@
However, recent trends have led to proposed cache coherent peripheral interconnects, and the extension of processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024). The more far-reaching are fully symmetric: devices are full participants in a distributed directory-based cache coherence protocol, as in CXL.mem 3.0 (CXL Consortium, 2024).
@@
All these workloads incur a significant latency penalty when using DMA for data transfer. Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads. For example, CPU execution and I/O were decoupled using descriptor rings because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
This is no longer true. Modern servers now have 100s of cores (Ampere Computing, 2024) and when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset (Marty et al., 2019). Most of the cores have nothing to do except handle small messages coming from the network.
@@
All this led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware. We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017; Liew et al., 2025; Schuh et al., 2024), but we push the question much further: given a future cache-coherent interconnect, what new ways of using PIO might be compelling for data center applications?
We are not proposing abandoning the DMA model, or some of its enhancements we survey in Section 6, for large-transfer throughput-oriented applications, which can keep riding the current hardware bandwidth curve. Instead, we argue it is time to examine complementary alternatives.
Figure 1. PCIe XDMA invocation latency comparison.
Figure 2. PCIe PIO invocation latency comparison.
@@ 3. Experimental platform
Both nodes have PCIe interfaces: the ThunderX-1 has a PCIe Gen.3 x8 interface, while the FPGA has a PCIe Gen.3 x16 interface. In this work we connect these two interfaces together with a cable; this replicates a PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
@@
We therefore calibrate both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 5.1. The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
@@
XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much less significant. Single transaction latency is almost constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
The performance difference between the machines is due to several factors. For small transfers, the CPU overhead of descriptor setup dominates, and an x86 core is simply faster than a ThunderX-1 core. In addition, the memory system is somewhat faster on the PC. The factor of 2 difference in bandwidth between the two platforms (PCIe Gen3 x8 vs. x16) does not appear to be a factor in these experiments.
@@
In the second calibration experiment the CPU performs PIO reads and writes over PCIe to the FPGA. PCIe is complex, and we spent considerable time optimizing CPU-initiated reads and writes for different platforms. We describe here how we arrived at optimally-efficient PIO latency to the FPGA over PCIe.
We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3) – this is basically uncached MMIO.
@@
Figure 3. Invoking with PIO & PCIe; error handling omitted.
@@
x86 has more optimization opportunities for PIO reads and writes. The Intel Data Streaming Accelerator (Intel, 2024) offers MOVDIR64B and ENQCMD instructions, allowing 64-byte PIO writes with write-combining mappings and user-space descriptor submission without driver intervention. Mapping memory with write-through attributes (as in Tide (Humphries et al., 2024)), allows cache line-sized PIO reads. These architecture-specific extensions are not available on ThunderX-1.
Figure 2 shows the best figures we could obtain on both platforms. For transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA. Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms). Writes, in contrast, are combined and pipelined at the PCIe interface.
Note this benchmark permits reordering of writes, and therefore already reflects the performance achievable using the reorder unit proposed in  (Liew et al., 2025).
@@ 4. PIO over a coherent interconnect
While PIO over PCIe outperforms DMA on our experimental hardware over a range of payload sizes (consistent with the findings of other authors), the main contribution of this paper to show the further benefits of using a cache-coherent interconnect. We present a family of message protocols, all of which run over an existing MESI-like coherence protocol, which substantially outperform conventional PIO and DMA.
Figure 4. FastForward-style coherent messaging
Fast CPU–CPU message passing
Our starting point is a set of software protocols used to communicate over shared memory between CPUs in a cache-coherent system, including FastForward (Giacomoni et al., 2008) and the NIC emulation used in CC-NIC (Schuh et al., 2024), and exemplified in Figure 4. These target machines with directory-based coherence which supports direct cache-to-cache forwarding of lines without writebacks, e.g. MOESI (Nagarajan et al., 2020).
In these protocols, a receiving core generally spins reading from a cache line (usually part of a larger array organized as a ring buffer), holding it in cache in Shared state. The sending core starts to write to the line, causing it to be fetched in Exclusive state and invalidated in the receiver’s cache. When the receiving core next polls the cache line, it is fetched from the sending core and arrives back in Shared state on the receiver, where it is immediately loaded into CPU registers.
In practice, the sender generally writes the entire line at once (exploiting a write buffer), so that in the fast case the line is transferred in two round-trips over the interconnect, as shown in Figure 4. Techniques like including a “finished” flag in the cache line data can be used to handle cases where the line is transferred to the receiver before it has been completely written.
These protocols have excellent performance, and on modern hardware achieve register-to-register latencies in the low hundreds of nanoseconds. They are easily parallelized to many cache lines (with a single polled flag for many lines of payload, plus the necessary barriers) to achieve excellent throughput. This basic technique is used, for example, CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket. However, they have several limitations.
The first is that the CPUs must busy-wait polling memory. This does not generate interconnect traffic since the line being polled is in local cache, but does consume energy. Moreover, optimal performance depends on timing, and if the receiver polls too frequently, the line can be forwarded multiple times before it is complete, leading to wasted interconnect round-trips and increased latency. For these reasons, these protocols are rarely used for real-world Inter-Process Communication (IPC), even when it makes sense to dedicate cores to individual clients and servers for long periods.
@@
The key insight behind this paper is that a device on a symmetric coherent interconnect sees more cache state and events than are visible to software, and can also initiate more operations than a CPU core can. Moreover, this does not require a cache on the device at all, indeed this is rarely appropriate: it is sufficient to generate and respond to individual coherence messages. We explain this in more detail.
First, the device now receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid). It is free to react internally to these events as it chooses, as long as it does not violate the protocol.
Second, the device can also issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
@@
This means requests to specific cache lines can be used to signal particular higher-level operations to the device, much as register access to traditional devices frequently has very different semantics to memory access.
(a) Write to device
(b) Read from device
(c) Read and write combined
@@
CPU-device message passing with coherence
We can now present a family of related point-to-point messaging protocols which can be built over a symmetric, directory-based cache coherence protocol where one end-point (the device) has direct access to the low-level cache messages.
Figure 5 shows three such protocols, all with the same broad structure, operating on a pair of cache lines which are homed at the device. This allows the device to precisely track states of the cache lines used and greatly simplifies the protocol. They vary according to whether data flows (a) from CPU to device, (b) from device to CPU, or (c) both ways.
Time is on the vertical axis. Under each end-point (CPU or device) are two columns, one for each cache line, giving its state over time. The notation E->M within a cell indicates the silent upgrade of a line from Exclusive to Modified, while the arrows between cells indicate a state changes corresponding to visible interconnect messages.
The first and second protocols implement uni-directional message passing between the CPU and device, and are the basis for the Ethernet NIC presented in Section 5.2. The third option combines the two by packing a request and response into one invocation of the protocol (i.e. 2 round-trips) and is best suited for send-and-receive interactions like RPC. We demonstrate this in Section 5.1.
Each protocol requires just two interconnect round trips, beginning with the two lines in a defined quiescent state. At each point one of the lines (which we refer to as $A$) serves to transfer information from CPU to device, and the other ($B$) from device to CPU. Over the course of one transaction both lines migrate from one cache to the other, ending in the same (caching protocol) state in which the other began. The two then swap roles for the next iteration.
There is no actual cache on the device: any payload is delivered directly into the device’s execution pipeline, or read from an output register, as appropriate. The following describes the execution of variant (c) in detail:
@@
    $B$ starts Exclusive in the CPU cache, and $A$ Invalid, with the (directory) states the opposite on the device. Software writes payload data from registers into $B$, which (silently) transitions to Modified.
@@
    To initiate the exchange, software reads from $A$, triggering a load shared message to the device. The device interprets this as a signal that $B$ contains fresh data.
    This coupling of independent line states is the first deviation from ‘standard’ coherence, and is impossible to achieve purely in software. It is nevertheless entirely invisible to the CPU.
@@
    The device now stalls the CPU by not responding to the load of $A$ until it has data to send. Instead, it requests $B$ in Exclusive to fetch the payload, leaving it Invalid in the CPU’s cache.
@@
    The device uses the data in $B$ to computes a result which it returns as a result payload when it finally responds to the CPU’s load from $A$.